# Netan Dogra

### v2

@@ 5.1. Integral mixed extensions
The extension class is $2$-torsion, as it is equal to the class obtained by pulling back under the hyperelliptic involution. We have a short exact sequence
$0\to\wedge^{2}T_{2}J\to I^{2}/I^{3}\to\mathrm{Sym}^{2}T_{2}J\to 0$
The image of $[I/I^{3}]$ in $H^{1}(K,\mathrm{Hom}(I/I^{2},\mathrm{Sym}^{2}T_{2}J))$ is equal to the class of $I\mathrm{Sym}^{\bullet}T_{2}J/I^{3}$, where $\mathrm{Sym}^{\bullet}T_{2}J$ is the symmetric algebra of $T_{2}J$ over $\mathbb{Z}_{2}$, viewed as a $\mathbb{Z}_{2}[\![\Pi]\!]$-module. This is extension class is trivial by the direct sum decomposition of the symmetric algebra as a Galois module. Hence it would be enough to show that $H^{1}(K,\mathrm{Hom}(T_{2}J,\wedge^{2}T_{2}J))[2]=0$.
This need not be true, but if we let $D\subset\mathbb{A}^{2g+1}_{K}$ denote the discriminant locus corresponding to inseparable monic degree $12g+$ 1 polynomials, then the class of $I/I^{3}$ in $H^{1}(K,\mathrm{Hom}(I/I^{2},I^{2}/I^{3}))$ is the specialisation of the class in $H^{1}(\pi_{1}^{\mathrm{\acute{e}t}}(\mathbb{A}^{2g+1}_{K}-D),\mathrm{Hom}(I/I^{2},I^{2}/I^{3}))$ corresponding to the natural hyperelliptic curve with a marked Weierstrass point over $\mathbb{A}^{2g+1}-D$. This class is also $2$-torsion, hence it is enough to show that $H^{1}(\pi_{1}^{\mathrm{\acute{e}t}}(\mathbb{A}^{2g+1}_{K}-D),\mathrm{Hom}(I/I^{2},\wedge^{2}T_{2}J))[2]=0$. This in turn reduces to showing that
$\mathrm{Hom}_{S_{2g+1}}(M,\wedge^{2}M)=0$
where $M$ is the standard $2g$-dimensional $\mathbb{F}_{2}$-representation of $S_{2g+1}$. This follows from the structure of Specht modules in positive characteristic [Jam78, §4]. ∎
We deduce from Lemma 21 that, for arbitrary $b,z\in X(K)$, the class of $I\mathbb{Z}_{2}[\![\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{K}};b,z)]\!]/I^{3}$ in $H^{1}(K,\mathrm{Hom}(I/I^{2},I^{2}/I^{3}))$ lies in the image of $H^{1}(K,T_{2}J)\times H^{1}(K,T_{2}J)$ under the natural map
$T_{2}J\times T_{2}J\to\mathrm{Hom}(I/I^{2},I^{2}/I^{3})$
given by left and right multiplication.
Suppose that $f$ is irreducible over $\mathbb{Q}_{2}$. Let $i_{f}:D_{f}\hookrightarrow H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$ under the right and left multiplication maps, and let $i:D\hookrightarrow H^{1}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$. Let $E_{f}(T_{2}J)$ be a crystalline extension
@@
defining an element of $H^{1}(\mathbb{Q}_{2},M_{1})$. If the class of $M_{1}$ in $\operatorname{Ext}^{1}(T_{2}J,I^{2}/I^{3})$ lies in the image of $D$ then we will denote by $[M]\in H^{1}(\mathbb{Q}_{2},E(T_{2}J))$ the image of this extension class under the map
@@
Since $c_{1}\cup c_{2}=\sum\pi_{1*}(c_{1})\cup\pi_{2*}(c_{2})=0$, there is an integral mixed extension $M$ of $c_{1}$ and $c_{2}$. Then $[M]-\sum[M_{i}]$ is an element of $H^{1}(\mathbb{Q}_{2},E(T_{2}J))$ which lies in the image of $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$, by the exact sequence
@@
In particular, we may twist $M$ by a class $\xi$ in $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$ to obtain a mixed extension $M^{\xi}$ with the property that $[M]=\sum[M_{i}]$ in $H^{1}(\mathbb{Q},E(T_{2}J))$. In particular, since all the $M_{i}$ are in the image of $H^{1}_{f}(\mathbb{Q}_{2},E_{f}(T_{2}J))$, so is $M^{\xi}$, and hence $M^{\xi}$ is crystalline by Lemma 22. ∎
We obtain the following criterion for liftability which is amenable to computations. First recall that for any $\mathbb{F}_{2}$-vector space $M$ we may view $M$ as a subspace of $M^{\otimes 2}$ via the homomorphism $m\mapsto m\otimes m$. Let $Y:=X-\{\infty\}$ and let
$\Theta_{1}:\mathbb{F}_{2}[Y(\mathbb{Z}_{2})\times Y(\mathbb{Z}_{2})]\to J(\mathbb{Q}_{2})/2J(\mathbb{Q}_{2})$
denote the homomorphism
$(P,Q)\mapsto[P]-[Q]$
Let
$\Theta_{2}:\mathrm{Ker}(\Theta_{1})\to(J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2})^{\otimes 2}/J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$
denote the homomorphism
$\sum(P_{i},Q_{i})\mapsto\sum_{i\neq j}(P_{i}-Q_{i})(P_{j}-Q_{j})+\sum_{i}(P_{i}-Q_{i})(Q_{i}-\{\infty\}).$
Let
$\Theta_{3}:\mathrm{Ker}(\Theta_{1})\to\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$
denote the homomorphism
$(P,Q)\mapsto(P-\{\infty\})(Q-\{\infty\})+(Q-\{\infty\})^{2}.$
@@
Suppose $\Theta_{1}$ and $\Theta_{2}\oplus\Theta_{3}$ are surjective. Then for every $b,z\in X(\mathbb{Q}_{2})$ such that $z-b\in 4J(\mathbb{Q}_{2})$, the lift of $j_{\Gamma,b}(z)$ to $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is in the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$.
@@
[…]
$\mathbb{Z}_{2}[Y(\mathbb{Z}_{2})\times Y(\mathbb{Z}_{2})]\to\widehat{J}\otimes_{\mathbb{Z}_{2}}\widehat{J}\times\mathrm{Sym}^{2}\widehat{J}$
$(P,Q)\mapsto((P-Q)\otimes(P-w(Q)),(P-Q)^{2})$
has image equal to the kernel $M$ of the map
$\widehat{J}\otimes_{\mathbb{Z}_{2}}\widehat{J}\times\mathrm{Sym}^{2}\widehat{J}\to\left(\widehat{J}\otimes_{\mathbb{Z}_{2}}\widehat{J}\otimes\mathbb{F}_{2}\times\mathrm{Sym}^{2}\widehat{J}\otimes\mathbb{F}_{2}\right)/J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2},$
where $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$ is embedded diagonally. Indeed it is clear that the image is contained in this subgroup, and to check it surjects onto it is enough to check it mod $2$ (i.e. that it surjects onto $M\otimes\mathbb{F}_{2}$). The vector space $M\otimes\mathbb{F}_{2}$ is an extension of $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$ by $(J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2})^{\otimes 2}/J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}\oplus\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$, and the map
$\mathbb{F}_{2}[Y(\mathbb{Z}_{2})\times Y(\mathbb{Z}_{2})]\to M\otimes\mathbb{F}_{2}$
is an extension of $\Theta_{1}$ by $\Theta_{2}\oplus\Theta_{3}$. Hence, given $P,Q$ such that $P-Q\in 4\cdot J(\mathbb{Q}_{2})$, there are $\lambda_{i}$ in $\mathbb{Z}_{2}$ and $P_{i},Q_{i}$ in $Y(\mathbb{Z}_{2})$ such that
$(P-Q)\otimes(P-w(Q))=\sum_{i}4\lambda_{i}\cdot(P_{i}-Q_{i})\otimes(P_{i}-w(Q_{i}))$
in $\widehat{J}\otimes\widehat{J}$. Hence $j_{\Gamma,b}(z)$ admits a crystalline lift by Lemma 23. ∎
@@ 6.1. Galois cohomology of number fields
For $X$ as above, $p$ a prime of good reduction and $m<2r-1$,
@@ 6.3. Dimension bounds via descent
$0\to\mathcal{O}_{L,S}^{\times}/\mathcal{O}_{L,S}^{\times p}\to(L^{\times}/L^{\times p})_{S}\to\mathrm{Cl}(\mathcal{O}_{L,S})[p]\to 0$
@@ 7.1. Computing $X(\mathbb{Q}_{2})_{2}$
Then the generalised height formalism, together with the verification of the rank part of the Bloch–Kato conjectures, tells us that $\Theta$ maps into an $N:=r(r+3)/2+(g-1)(g+2)/2$-dimensional subspace of the target, which is an $N+g^{2}-g$ dimensional vector space.
@@
This is not yet an identity which can be tested on $p$-adic points, because it involves the image of $z-b$ in $J(\mathbb{Q})\otimes\mathbb{Q}_{p}$, which does not inject into $H^{0}(X_{\mathbb{Q}_{p}},\mathbb{Q}_{p})^{*}$. We can remedy this by elimination of variables.
@@ 8. $2$-adic Coleman integration
where $f(x)=\sum_{i=0}^{2g+1}a_{i}x^{i}$ has degree $12g+$ 1. Let $Q(x,y):=y^{2}-y-f(x)$. Define $(y_{n})$ by
@@
This implies that $2^{n}$ divides $Q(x^{2},y_{n})$, and hence implies the first part of the lemma. We may write $y_{n}=f_{1}(x)+yf_{2}(x)$ for unique $f_{1},f_{2}\in\mathbb{Z}_{2}[x]$. Then [DV06, Lemma 1] implies that the degree of $f_{1}$ and $f_{2}$ is at most $26(n-1)+$ 2, which implies part (2) of the Lemma. ∎
@@
where $g_{i},h_{i}\in\mathbb{Z}[x]$ have degree less than $d+i-1$. An explicit convergence estimate is given in [DV02, Lemma 1].
@@ 8.1. The universal unipotent connection on $X$
[…]
@@ 8.1.1. Frobenius structure
[…]
@@ 8.1.3. Extending to infinity
[…]
@@ 8.2. Computing the generalised height
$0\to(H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*\otimes 2}/\mathbb{Q}_{p}(1))\otimes_{\mathbb{Q}_{p}}\mathcal{O}_{X}\to\mathcal{E}_{2}\to\mathcal{E}_{1}\to 0$
@@
For each $x\in X(\mathbb{Q}_{p})$, $x^{*}\mathcal{E}$ then has the structure of a mixed extension of filtered $\phi$-modules with graded pieces $\mathbb{Q}_{p},H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}$ and $\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}/\mathbb{Q}_{p}(1)$.
@@ 9.1. The BMSST curve
where $\mathrm{Im}(U)$ is the image of $U$ in $U^{\prime}:=\mathrm{Ker}(L_{\mathfrak{q}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K_{\mathfrak{p}}^{\times}\otimes\mathbb{F}_{2})$. To do this, we need to compute the crystalline subspace. We do this using our results on integral crystalline lifts and nonabelian $(x-T)$ maps. More precisely, we first computationally verify that $X$ satisfies the conditions of Lemma 24. Hence, if $b$ and $z$ are $\mathbb{Z}_{2}$-points of $X-\{\infty\}$ such that $z-b\in 4\cdot J(\mathbb{Q}_{2})$, then Proposition 2 defines an element of $U^{\prime}$ which is a lift of the class $j_{\Gamma,b}(z)$, and Lemma 24 implies that this lies in $\mathcal{L}$ if the image of $X(\mathbb{Q}_{2})$ in $\mathrm{Sym}^{2}H^{1}(\mathbb{Q}_{2},J[2])$ spans $\mathrm{Sym}^{2}H^{1}_{f}(\mathbb{Q}_{2},J[2])$. This reduces verification of (22) to calculating a basis of $\mathcal{O}_{L}^{\times}$, and calculating whether an element of $L_{\mathfrak{q}}$ is a square. Both of these calculations have been studied in previous work on implementing the Chabauty–Coleman method for hyperelliptic curves (see [Sto01]). We carried out this calculation in magma, searching over $\mathbb{Z}_{2}$ points of $X-\{\infty\}$ to determine a basis of $\mathcal{L}$ and verify (22).
@@
and $-54h(P_{1})-6h(P_{2})-6h(P_{3})+3h(P_{5})+2h(P_{6})+h(P_{8})$ span a two dimensional subspace of the four dimensional vector space

### v3

@@ 5.1. Integral mixed extensions
One may choose a $\mathbb{Z}_{p}$-module splitting $\alpha:T_{0}\oplus T_{1}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}M/T_{2}\simeq T_{0}\oplus T_{1}$ with respect to which the action of $G_{\mathbb{Q}_{p}}$ is given by
$\left(\begin{array}[]{cc}\rho_{0}&0\\ p\cdot c(g)&\rho_{1}\end{array}\right).$
Then $\alpha(T_{0}\oplus pT_{1})$ is a Galois stable sublattice. Taking the preimage of $\alpha(T_{0}\oplus pT_{1})$ in $M$ gives the lemma. ∎
We now turn to the problem of describing the structure of $E_{2}(b)$. By definition this is a direct sum of $\mathbb{Z}_{2}$ and $IE_{2}(b)$. Hence the problem is to describe the structure of $I\mathbb{Z}_{2}[\![\Pi_{2}]\!]/I^{3}$ as an extension of $T_{2}J$ by $T_{2}J^{\otimes 2}$. This is closely related to the class of the Ceresa cycle of $(X,b)$ in Galois cohomology, by work of Hain and Matsumoto [HM05]. The relation integrally (rather than over $\mathbb{Q}_{2}$) has some additional subtleties which were recently studied by Bisogno, Li, Litt and Srinivasan [BLLS23]. They show that, if $\infty$ denotes a rational Weierstrass point, then the class of $I\mathbb{Z}_{2}[\![\Pi_{2}(\infty)]\!]/I^{3}$ is $2$-torsion.
Let $\theta:H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\times H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\to H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,T_{J}^{\otimes 2}))$ be the map induced by the left and right multiplication maps
(16) $T_{2}J\times T_{2}J\to\mathrm{Hom}(I/I^{2},I^{2}/I^{3}).$
We deduce that, for arbitrary $b,z\in(X-\infty)(\mathbb{Z}_{2})$, the class of $IE_{2}(b,z)$ in $H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,T_{J}^{\otimes 2}))$ is equal to
(17) $E_{2}(\infty)]+\theta(z-\infty,b-\infty).$
In particular, this implies the following.
Lemma 22.
For any $b,z\in(X-\infty)(\mathbb{Z}_{2})$, we have
$2[IE_{2}(b,z)]\in 2\cdot\mathrm{Im}(\theta).$
Suppose that $f$ is irreducible over $\mathbb{Q}_{2}$. Let $i_{f}:D_{f}\hookrightarrow H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $\theta$, and let $i:D\hookrightarrow H^{1}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$. Let $E_{f}(T_{2}J)$ be a crystalline extension
@@
defining an element of $H^{1}(\mathbb{Q}_{2},M_{1})$. If the class of $M_{1}$ in $\operatorname{Ext}^{1}(T_{2}J,I^{2}/I^{3})$ lies in the image of $D$ then we will denote by $M]\in H^{1}(\mathbb{Q}_{2},E(T_{2}J))$ the image of this extension class under the map
@@
Since $c_{1}\cup c_{2}=\sum\pi_{1*}(c_{1})\cup\pi_{2*}(c_{2})=0$, there is an integral mixed extension $M$ of $c_{1}$ and $c_{2}$. Then $M]-\sum[M_{i}]$ is an element of $H^{1}(\mathbb{Q}_{2},E(T_{2}J))$ which lies in the image of $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$, by the exact sequence
@@
In particular, we may twist $M$ by a class $\xi$ in $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$ to obtain a mixed extension $M^{\xi}$ with the property that $M]=\sum[M_{i}]$ in $H^{1}(\mathbb{Q},E(T_{2}J))$. In particular, since all the $M_{i}$ are in the image of $H^{1}_{f}(\mathbb{Q}_{2},E_{f}(T_{2}J))$, so is $M^{\xi}$, and hence $M^{\xi}$ is crystalline by Lemma 23. ∎
We obtain the following criterion for liftability which is amenable to computations. Let $\widehat{J}$ denote the $2$-adic completion of $J(\mathbb{Q}_{2})$.
@@
Suppose the homomorphism of $\mathbb{Z}_{2}$-modules
$\mathbb{Z}_{2}[(X-\infty)(\mathbb{Z}_{2})\times(X-\infty)(\mathbb{Z}_{2})]\to\widehat{J}^{\otimes 2}\times\widehat{J}^{\otimes 2}$
sending $(z_{1},z_{2})$ to $((z_{1}-z_{2})\otimes(z_{1}-\infty),(z_{1}-z_{2})\otimes(z_{2}-\infty))$ has image equal to the preimage of $\mathrm{Sym}^{2}\widehat{J}\subset\widehat{J}^{\otimes 2}$ under the map
$\widehat{J}^{\otimes}2\to\widehat{J}^{\otimes 2}$
sending $(v,w)$ to vwv-w. Then for ever $b,z\in(X-\infty)(\mathbb{Z}_{2})$ such that $z-b\in 4J(\mathbb{Q}_{2})$, the lift of $j_{\Gamma,b}(z)$ to $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is in the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$.
@@
By (16) and Lemma 24, the conditions imply that there is a crystalline mixed extension $M$ who image in $H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)\otimes H^{1}_{f}(\mathbb{Q}_{2},T_{2}J^{*}\otimes T_{2}J^{\otimes 2})$ is equal to twice that of $E_{2}(b,z)$. Rescaling as in Lemma 21 we may replace $E_{2}(b,z)$ with a mixed extension $M^{\prime}\subset E_{2}(b,z)$ whose class in $H^{1}_{f}(\mathbb{Q}_{2},T_{2}J^{*}\otimes I^{2}/I^{3})$ lies in the image of $\theta$. Hence taking $M^{\prime}]-2[M]$ gives an extension of $\mathbb{Z}_{2}$ by $T_{2}J^{\otimes 2}$ realising the extension class. ∎
@@ 6.1. Galois cohomology of number fields
For $X$ as above, $p$ a prime of good reduction and m<2r1m<2r-1,
@@ 6.3. Dimension bounds via descent
$0\to\mathcal{O}_{L,S}^{\times}/\mathcal{O}_{L,S}^{\times p}\to(L^{\times}/L^{\times p})_{S}\to\mathrm{Cl}(\mathcal{O}_{L,S})[p]\to 0.$
@@ 7.1. Computing $X(\mathbb{Q}_{2})_{2}$
Then the generalised height formalism, together with the verification of the rank part of the Bloch–Kato conjectures, tells us that $\Theta$ maps into an N:=r(r+3)/2+(g1)(g+2)/2N:=r(r+3)/2+(g-1)(g+2)/2-dimensional subspace of the target, which is an $N+g^{2}-g$ dimensional vector space.
@@
This is not yet an identity which can be tested on $p$-adic points, because it involves the image of zbz-b in $J(\mathbb{Q})\otimes\mathbb{Q}_{p}$, which does not inject into $H^{0}(X_{\mathbb{Q}_{p}},\mathbb{Q}_{p})^{*}$. We can remedy this by elimination of variables.
@@ 8. $2$-adic Coleman integration
where $f(x)=\sum_{i=0}^{2g+1}a_{i}x^{i}$ has degree $2g+1$. Let $Q(x,y):=y^{2}-y-f(x)$. Define $(y_{n})$ by
@@
This implies that $2^{n}$ divides $Q(x^{2},y_{n})$, and hence implies the first part of the lemma. We may write $y_{n}=f_{1}(x)+yf_{2}(x)$ for unique $f_{1},f_{2}\in\mathbb{Z}_{2}[x]$. Then [DV06, Lemma 1] implies that the degree of $f_{1}$ and $f_{2}$ is at most 6(n1)+26(n-1)+2, which implies part (2) of the Lemma. ∎
@@
where $g_{i},h_{i}\in\mathbb{Z}[x]$ have degree less than d+i1d+i-1. An explicit convergence estimate is given in [DV02, Lemma 1].
@@ 8.1. The universal unipotent connection on $X$
For any $z\in X(\mathbb{F}_{2})$, the connection $\mathcal{E}_{n}$ admits a rigid analytic trivialisation upon restriction to the residue disk ]z[]z[, i.e. the map
@@ 8.1.1. Frobenius structure
Given a lift $\phi$ of Frobenius on ]U[]U[, for all $x\in U(\mathbb{F}_{2})$ there is a unique point $\iota(x)\in]U[(\mathbb{Q}_{2})$ fixed by $\phi$, by Hensel’s lemma. Given an isocrystal $\mathcal{E}$ with Frobenius structure $F$, and a point $\overline{z}\in U(\mathbb{F}_{2})$ we hence obtain an automorphism $z_{0}^{*}(F)$ of $z_{0}^{*}\mathcal{E}$, where $z_{0}:=\iota(\overline{z})$. For general $z$ in ]U[]U[, we define $z^{*}F$ to be the automorphism $I_{z_{0},z}\circ z_{0}^{*}F\circ I_{z,z_{0}}$.
@@ 8.1.3. Extending to infinity
For our final calculations it will be important to compute the generalised height, and hence the matrix of Frobenius, at the residue disk at infinity. For this, we use overconvergence of our lift of Frobenius. Namely, if our Frobenius lift is defined on a strict neighbourhood $U$ of ]Y[]Y[ for an open subscheme $Y\subset X_{\mathbb{F}_{p}}$ (assume for simplicity that all points of XYX-Y are defined over $\mathbb{F}_{p}$), we may apply (21) to deduce that for any $z_{1},z_{2}\in]x[(\overline{\mathbb{Q}}_{p})$, we have
@@ 8.2. Computing the generalised height
$0\to(H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*\otimes 2}/\mathbb{Q}_{p}(1))\otimes_{\mathbb{Q}_{p}}\mathcal{O}_{X}\to\mathcal{E}_{2}\to\mathcal{E}_{1}\to 0.$
@@
For each $\in X(\mathbb{Q}_{p}), xx^{*}\mathcal{E}$ then has the structure of a mixed extension of filtered $\phi$-modules with graded pieces $\mathbb{Q}_{p},H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}$ and $\wedge^{2}H^{1}_{\mathrm{dR}}(X/\mathbb{Q}_{p})^{*}/\mathbb{Q}_{p}(1)$.
@@ 9.1. The BMSST curve
where $\mathrm{Im}(U)$ is the image of $U$ in $U^{\prime}:=\mathrm{Ker}(L_{\mathfrak{q}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K_{\mathfrak{p}}^{\times}\otimes\mathbb{F}_{2})$. To do this, we need to compute the crystalline subspace. We do this using our results on integral crystalline lifts and nonabelian (xT)(x-T) maps. More precisely, we first computationally verify that $X$ satisfies the conditions of Lemma 25. Hence, if $b$ and $z$ are $\mathbb{Z}_{2}$-points of $X-\{\infty\}$ such that $z-b\in 4\cdot J(\mathbb{Q}_{2})$, then Proposition 2 defines an element of $U^{\prime}$ which is a lift of the class $j_{\Gamma,b}(z)$, and Lemma 25 implies that this lies in $\mathcal{L}$ if the image of $X(\mathbb{Q}_{2})$ in $\mathrm{Sym}^{2}H^{1}(\mathbb{Q}_{2},J[2])$ spans $\mathrm{Sym}^{2}H^{1}_{f}(\mathbb{Q}_{2},J[2])$. This reduces verification of (24) to calculating a basis of $\mathcal{O}_{L}^{\times}$, and calculating whether an element of $L_{\mathfrak{q}}$ is a square. Both of these calculations have been studied in previous work on implementing the Chabauty–Coleman method for hyperelliptic curves (see [Sto01]). We carried out this calculation in magma, searching over $\mathbb{Z}_{2}$ points of $X-\{\infty\}$ to determine a basis of $\mathcal{L}$ and verify (24).
@@
and $54h(P_{1})-6h(P_{2})-6h(P_{3})+3h(P_{5})+2h(P_{6})+h(P_{8})$ span a two dimensional subspace of the four dimensional vector space